# From Timestamps to Versions: Version AoI in Single- and Multi-Hop Networks

### v1

@@ Appendix A Optimal On-Off Policy for the CMDP Problem
For a finite state space $S$, the growth condition in [35, Eq. 11.21] holds. Since the transition cost $C(s(t),a(t),s(t+1))\geq 0$ is bounded below, the conditions of [35, Corollary 12.2] are satisfied, ensuring the optimal solutions of the dual and primal problems coincide. Thus, the optimal solution to the primal cmdp (16) is found by solving $\sup_{\lambda\geq 0}\mathcal{L}(\lambda,\pi_{\lambda}^{\ast})$, where $\pi_{\lambda}^{\ast}$ comes from (20). Specifically, the optimal policy is obtained by first solving the unconstrained mdp (20) for fixed $\lambda$ to get $\pi_{\lambda}^{\ast}$, and then optimizing $\lambda$ as in (18). We proceed to prove that $\pi_{\lambda}^{\ast}$ is a threshold policy.
@@
$J^{\ast}\!\!+\!\mathbb{V}(s)\!=\!\!\!\min_{a\in\left\{0,1\right\}}\!{\!Q_{\lambda}(s,a)},\ \ \pi^{\ast}(s)\!\in\!\operatorname*{arg\,min}_{a\in\left\{0,1\right\}}{Q_{\lambda}(s,a)},$ (22)
@@
This implies a threshold policy: if $\text{d}\mathbb{V}(\Delta_{T})<0$ for some $\Delta_{T}$, then $\text{d}\mathbb{V}(\Delta)<0$ for all $\Delta\geq\Delta_{T}$, so the optimal action remains 111 for all such states. Using (17), we obtain:
@@
We prove by induction that $\mathbb{V}_{k}(\Delta\!^{-})\!\leq\!\mathbb{V}_{k}(\Delta\!^{+})$ for all $k\!\geq\!0$. For $k\!=\!0$, $\mathbb{V}_{0}(\Delta)\!=\!0$, so the claim holds. Assume $\mathbb{V}_{k}(\Delta\!^{-})\!\leq\!\mathbb{V}_{k}(\Delta\!^{+})$ and show $\mathbb{V}_{k+1}(\Delta\!^{-})\!\leq\!\mathbb{V}_{k+1}(\Delta\!^{+})$. Since $\mathbb{V}_{k+1}(\Delta)=\min\{\mathbb{V}_{k+1}^{0}(\Delta),\mathbb{V}_{k+1}^{1}(\Delta)\}$ with $\mathbb{V}_{k+1}^{0}\!=\!Q_{\lambda,k}(\Delta,0)$ and $\mathbb{V}_{k+1}^{1}\!=\!Q_{\lambda,k}(\Delta,1)$, it suffices to prove $\mathbb{V}_{k+1}^{0}(\Delta\!^{-})\!\leq\!\mathbb{V}_{k+1}^{0}(\Delta\!^{+})$ and $\mathbb{V}_{k+1}^{1}(\Delta\!^{-})\!\leq\!\mathbb{V}_{k+1}^{1}(\Delta\!^{+})$, since then $\min\{\mathbb{V}_{k+1}^{0}(\Delta\!^{-}),\mathbb{V}_{k+1}^{1}(\Delta\!^{-})\}\!\leq\!\min\{\mathbb{V}_{k+1}^{0}(\Delta\!^{+}),\mathbb{V}_{k+1}^{1}(\Delta\!^{+})\}$. By the induction hypothesis, all bracketed terms below are non-positive:
$\mathbb{V}$ (${}^{0}_{\!k+1}(\Delta\!^{-})\!-\!\mathbb{V}^{0}_{\!k+1}(\Delta\!^{+})\!=\!\big{[}\Delta\!^{-}\!-\!\Delta\!^{+}\big{]}$
@@
$\mathbb{V}$ (${}^{1}_{\!k+1}(\Delta\!^{-})\!-\!\mathbb{V}^{1}_{\!k+1}(\Delta\!^{+})\!=\!\bar{p}_{s}\big{[}\Delta\!^{-}\!\!-\!\Delta\!^{+}\big{]}$
@@
At each time slot, a transmission attempt occurs w.p. $\alpha$ and succeeds w.p. $p_{s}$, giving a success probability of $\alpha p_{s}$. Upon success, the VAoI resets to 0 or 111 depending on whether a new version is generated in that slot. Otherwise, if no transmission occurs or it fails (w.p. $1-\alpha p_{s}$), the VAoI increases by 0 or 111, again depending on version generation. The Markov chain with transition probabilities is shown in Fig. 12, and solving its balance equations as described below yields Proposition 1.
@@
[…]
@@
For each sampled Markov chain $Y^{(q)}(k)=\Delta(kD\!+\!q)$, the VAoI transitions from $Y^{(q)}(k)$ to $Y^{(q)}(k\!+\!1)$ depending on whether the transmission at $t\!=\!(k\!+\!1)D$ succeeds (or not) and on the number of source versions generated in the $q$ (or $D$) slots.
  • Successful transmission (w.p. $p_{s}$): the VAoI resets to zero and subsequently increases with the versions generated within $q$ time slots: it equals 0 if none are generated (w.p. $p_{s}b^{q}_{0}$); 111 if one is generated (w.p. $p_{s}b^{q}_{1}$); and so on, up to $n\leq q$, where $n$ versions are generated (w.p. $p_{s}b^{q}_{n}$). Version generation per slot follows i.i.d. Bernoulli with parameter $p_{g}$; thus, generating $z$ versions in $q$ slots follows $Bin(q,p_{g})$ w.p. $b^{q}_{z}\!=\!\binom{q}{z}p_{g}^{z}(1-p_{g})^{q-z}$.
  • Unsuccessful transmission (w.p. $1-p_{s}$): the VAoI increases by the number of versions generated during $D$ slots: 0 w.p. $(1-p_{s})b^{D}_{0}$, 111 w.p. $(1-p_{s})b^{D}_{1}$, and so on, up to $D$ w.p. $(1-p_{s})b^{D}_{D}$.
Fig. 13 illustrates the Markov chain states and their transitions to target states—0, 111, $1\!<\!n\!\leq\!q$, $q\!<\!n\!\leq\!D$, and $n\!>\!D$—together with the associated transition probabilities. By formulating the balance equations for each state, we derive:
@@
For $\Delta_{T}=0$, when the VAoI is zero, transmission decisions do not affect state transitions, since the VAoI depends solely on whether a new version is generated. Hence, the system dynamics and steady-state distribution are identical for $\Delta_{T}=0$ and $\Delta_{T}=1$. The case $\Delta_{T}=0$ represents an always-update policy, equivalent to a randomized stationary policy with $\alpha=1$. For $\Delta_{T}\geq 2$, if $\Delta<\Delta_{T}$, no transmission occurs; the VAoI increments by 111 if a new version is generated, otherwise it remains the same. If $\Delta\geq\Delta_{T}$, transmission occurs: on success, VAoI resets to 111 or 0 depending on whether a new version is generated; on failure, it increases by 111 if a new version is generated, or remains unchanged otherwise. The resulting Markov chain is depicted in Fig. 14, with balance equations provided for $\Delta_{T}\geq 2$:
@@
This implies that $\Delta_{T}$ must exceed a certain lower bound. Meanwhile, Lemma 2 shows that the average VAoI under the threshold policy (8) is an increasing function of $\Delta_{T}$, since it can be rewritten as: $\bar{\Delta}\!=\!\frac{\Delta_{T}}{2}\left(1\!-\!\frac{\beta}{(\Delta_{T}\!-\!1)p_{s}\!+\!\beta}\right)\!+\!\frac{p_{g}}{p_{s}}.$ The constant term $\frac{p_{g}}{p_{s}}$ remains fixed; the first term increases with $\Delta_{T}$ due to the linear growth of $\frac{\Delta_{T}}{2}$ and the rising value of $\left(1\!-\!\frac{\beta}{(\Delta_{T}-1)p_{s}+\beta}\right)$, since $\frac{\beta}{(\Delta_{T}-1)p_{s}+\beta}$ decreases as $\Delta_{T}$ grows. Therefore, minimizing the average VAoI under the update rate constraint entails selecting the smallest integer $\Delta_{T}$ that satisfies (25). However, exact equality may not always be attainable, as $\Delta_{T}$ must be integer-valued. To achieve $R(\Delta_{T})\!=\!\alpha$, a randomized mixture policy can be employed [37][35, Sec. 6.3], combining two thresholds $\Delta_{T}^{\ast}$ and $\Delta_{T}^{\ast}\!-\!1$, where $R(\Delta_{T}^{\ast})\!\leq\!\alpha$ and $R(\Delta_{T}^{\ast}\!-\!1)\!>\!\alpha$, resulting in (9). The threshold $\Delta_{T}^{\ast}$ is applied w.p. $\gamma$ and $\Delta_{T}^{\ast}\!-\!1$ with $1\!-\!\gamma$, where $\gamma R(\Delta_{T}^{\ast})+(1-\gamma)R(\Delta_{T}^{\ast}-1)=\alpha$, leading to (10). This mixed threshold policy constitutes the optimal solution to the cmdp under the average rate constraint (1) [37]. ∎
@@
where $\eta_{1}=V_{S}(t)-V_{S}(t-1)$ represents the number of new versions generated by the source in the most recent slot (either 0 or 111). If the latest transmission fails (w.p. $1-\rho_{i}$), but the previous one at $t-1$ succeeds (w.p. $\rho_{i}$), then $V_{i+1}(t)=V_{i}(t-2)$, and the VAoI becomes:
@@
=$\!\overset{(a)}{=}\!{\sum_{\tau=N}^{\infty}\!\mathbb{P}(\tau_{N}\!=\!\tau)\mathbb{E}\left[\Delta_{1}(t\!-\!\tau)\right]}\!+\!p_{g}\!\sum_{i=1}^{N}\frac{1}{\rho_{i}},$
@@
Equality $(b)$ follows directly from the relation $\lim_{t\rightarrow\infty}\mathbb{E}[\Delta_{1}(t-\tau)]=\lim_{t\rightarrow\infty}\mathbb{E}[\Delta_{1}(t)]$, which holds for an ergodic and integrable dtmc $\Delta_{1}(t)$, where $\mathbb{E}\big{[}|\Delta_{1}(t)|\big{]}<\infty$ [34, Sec. 1.10]. This condition is satisfied by the dtmcs presented in Section IV under all three policies, for which the steady-state distribution is stationary and therefore time-invariant. ∎

### v2

@@ Appendix A Optimal On-Off Policy for the CMDP Problem
For a finite state space $S$, the growth condition in [4, Eq. 11.21] holds. Since the transition cost $C(s(t),a(t),s(t+1))\geq 0$ is bounded below, the conditions of [4, Corollary 12.2] are satisfied, ensuring the optimal solutions of the dual and primal problems coincide. Thus, the optimal solution to the primal cmdp (16) is found by solving $\sup_{\lambda\geq 0}g(\lambda)$, where $\pi_{\lambda}^{\ast}$ comes from (20). Specifically, the optimal policy is obtained by first solving the unconstrained mdp (20) for fixed $\lambda$ to get $\pi_{\lambda}^{\ast}$, and then optimizing $\lambda$ as in (18). We proceed to prove that $\pi_{\lambda}^{\ast}$ is a threshold policy.
@@
$J_{\lambda}^{*}\!\!+\!\mathbb{V}(s)\!=\!\!\!\min_{a\in\left\{0,1\right\}}\!{\!Q_{\lambda}(s,a)},\ \ \pi^{\ast}(s)\!\in\!\operatorname*{arg\,min}_{a\in\left\{0,1\right\}}{Q_{\lambda}(s,a)},$ (22)
@@
This implies a threshold policy: if $\text{d}\mathbb{V}(\Delta_{T})<0$ for some $\Delta_{T}$, then $\text{d}\mathbb{V}(\Delta)<0$ for all $\Delta\geq\Delta_{T}$, so the optimal action remains 11 for all such states. Using (17), we obtain:
@@
We prove by induction that $\mathbb{V}_{k}(\Delta\!^{-})\!\leq\!\mathbb{V}_{k}(\Delta\!^{+})$ for all $k\!\geq\!0$. For k=0k\!=\!0, $\mathbb{V}_{0}(\Delta)\!=\!0$, so the claim holds. Assume $\mathbb{V}_{k}(\Delta\!^{-})\!\leq\!\mathbb{V}_{k}(\Delta\!^{+})$ and show $\mathbb{V}_{k+1}(\Delta\!^{-})\!\leq\!\mathbb{V}_{k+1}(\Delta\!^{+})$. Since $\mathbb{V}_{k+1}(\Delta)=\min\{\mathbb{V}_{k+1}^{0}(\Delta),\mathbb{V}_{k+1}^{1}(\Delta)\}$ with $\mathbb{V}_{k+1}^{0}\!=\!Q_{\lambda,k}(\Delta,0)$ and $\mathbb{V}_{k+1}^{1}\!=\!Q_{\lambda,k}(\Delta,1)$, it suffices to prove $\mathbb{V}_{k+1}^{0}(\Delta\!^{-})\!\leq\!\mathbb{V}_{k+1}^{0}(\Delta\!^{+})$ and $\mathbb{V}_{k+1}^{1}(\Delta\!^{-})\!\leq\!\mathbb{V}_{k+1}^{1}(\Delta\!^{+})$, since then $\min\{\mathbb{V}_{k+1}^{0}(\Delta\!^{-}),\mathbb{V}_{k+1}^{1}(\Delta\!^{-})\}\!\leq\!\min\{\mathbb{V}_{k+1}^{0}(\Delta\!^{+}),\mathbb{V}_{k+1}^{1}(\Delta\!^{+})\}$. By the induction hypothesis, all bracketed terms below are non-positive:
$\mathbb{V}$ $^{0}_{\!k+1}(\Delta\!^{-})\!-\!\mathbb{V}^{0}_{\!k+1}(\Delta\!^{+})\!=\!\big[\Delta\!^{-}\!-\!\Delta\!^{+}\big]$
@@
$\mathbb{V}$ $^{1}_{\!k+1}(\Delta\!^{-})\!-\!\mathbb{V}^{1}_{\!k+1}(\Delta\!^{+})\!=\!\bar{p}_{s}\big[\Delta\!^{-}\!\!-\!\Delta\!^{+}\big]$
@@
At each time slot, a transmission attempt occurs w.p. $\alpha$ and succeeds w.p. $p_{s}$, giving a success probability of $\alpha p_{s}$. Upon success, the VAoI resets to 0 or $1$ depending on whether a new version is generated in that slot. Otherwise, if no transmission occurs or it fails (w.p. $1-\alpha p_{s}$), the VAoI increases by 0 or $1$, again depending on version generation. The Markov chain with transition probabilities is shown in Fig. 12, and solving its balance equations as described below yields Proposition 1.
@@
Figure 13: State transitions to target states (a) 0(a)\ 0, (b) 1(b)\ 1, $(c)\ 1\!<\!n\!\leq\!q$, $(d)\ q\!<\!n\!\leq\!D$, and (e)n>D(e)\ n\!>\!D in the Markov chain under a uniform policy.
@@
For each sampled Markov chain $Y^{(q)}(k)=\Delta(kD\!+\!q)$, the VAoI transitions from $Y^{(q)}(k)$ to $Y^{(q)}(k\!+\!1)$ depending on whether the transmission at t=(k+1)Dt\!=\!(k\!+\!1)D succeeds (or not) and on the number of source versions generated in the $q$ (or $D$) slots.
  • Successful transmission (w.p. $p_{s}$): the VAoI resets to zero and subsequently increases with the versions generated within $q$ time slots: it equals 0 if none are generated (w.p. $p_{s}b^{q}_{0}$); $1$ if one is generated (w.p. $p_{s}b^{q}_{1}$); and so on, up to $n\leq q$, where $n$ versions are generated (w.p. $p_{s}b^{q}_{n}$). Version generation per slot follows i.i.d. Bernoulli with parameter $p_{g}$; thus, generating $z$ versions in $q$ slots follows $Bin(q,p_{g})$ w.p. $b^{q}_{z}\!=\!\binom{q}{z}p_{g}^{z}(1-p_{g})^{q-z}$.
  • Unsuccessful transmission (w.p. $1-p_{s}$): the VAoI increases by the number of versions generated during $D$ slots: 0 w.p. $(1-p_{s})b^{D}_{0}$, $1$ w.p. $(1-p_{s})b^{D}_{1}$, and so on, up to $D$ w.p. $(1-p_{s})b^{D}_{D}$.
Fig. 13 illustrates the Markov chain states and their transitions to target states—0, 11, $1\!<\!n\!\leq\!q$, $q\!<\!n\!\leq\!D$, and n>Dn\!>\!D—together with the associated transition probabilities. By formulating the balance equations for each state, we derive:
@@
For $\Delta_{T}=0$, when the VAoI is zero, transmission decisions do not affect state transitions, since the VAoI depends solely on whether a new version is generated. Hence, the system dynamics and steady-state distribution are identical for $\Delta_{T}=0$ and $\Delta_{T}=1$. The case $\Delta_{T}=0$ represents an always-update policy, equivalent to a randomized stationary policy with $\alpha=1$. For $\Delta_{T}\geq 2$, if $\Delta<\Delta_{T}$, no transmission occurs; the VAoI increments by $1$ if a new version is generated, otherwise it remains the same. If $\Delta\geq\Delta_{T}$, transmission occurs: on success, VAoI resets to $1$ or 0 depending on whether a new version is generated; on failure, it increases by $1$ if a new version is generated, or remains unchanged otherwise. The resulting Markov chain is depicted in Fig. 14, with balance equations provided for $\Delta_{T}\geq 2$:
@@
This implies that $\Delta_{T}$ must exceed a certain lower bound. Meanwhile, Lemma 2 shows that the average VAoI under the threshold policy (8) is an increasing function of $\Delta_{T}$, since it can be rewritten as: $\bar{\Delta}\!=\!\frac{\Delta_{T}}{2}\left(1\!-\!\frac{\beta}{(\Delta_{T}\!-\!1)p_{s}\!+\!\beta}\right)\!+\!\frac{p_{g}}{p_{s}}.$ The constant term $\frac{p_{g}}{p_{s}}$ remains fixed; the first term increases with $\Delta_{T}$ due to the linear growth of $\frac{\Delta_{T}}{2}$ and the rising value of $\left(1\!-\!\frac{\beta}{(\Delta_{T}-1)p_{s}+\beta}\right)$, since $\frac{\beta}{(\Delta_{T}-1)p_{s}+\beta}$ decreases as $\Delta_{T}$ grows. Therefore, minimizing the average VAoI under the update rate constraint entails selecting the smallest integer $\Delta_{T}$ that satisfies (25). However, exact equality may not always be attainable, as $\Delta_{T}$ must be integer-valued. To achieve $R(\Delta_{T})\!=\!\alpha$, a randomized mixture policy can be employed [9][4, Sec. 6.3], combining two thresholds $\Delta_{T}^{\ast}$ and $\Delta_{T}^{\ast}\!-\!1$, where $R(\Delta_{T}^{\ast})\!\leq\!\alpha$ and $R(\Delta_{T}^{\ast}\!-\!1)\!>\!\alpha$, resulting in (9). The threshold $\Delta_{T}^{\ast}$ is applied w.p. $\gamma$ and $\Delta_{T}^{\ast}\!-\!1$ with $1\!-\!\gamma$, where $\gamma R(\Delta_{T}^{\ast})+(1-\gamma)R(\Delta_{T}^{\ast}-1)=\alpha$, leading to (10). This mixed threshold policy constitutes the optimal solution to the cmdp under the average rate constraint (1) [9]. ∎
@@
where $\eta_{1}=V_{S}(t)-V_{S}(t-1)$ represents the number of new versions generated by the source in the most recent slot (either 0 or $1$). If the latest transmission fails (w.p. $1-\rho_{i}$), but the previous one at t1t-1 succeeds (w.p. $\rho_{i}$), then $V_{i+1}(t)=V_{i}(t-2)$, and the VAoI becomes:
@@
$\begin{cases}\!\Delta_{2}(t)\!=\!\Delta_{1}(t\!-\!m_{1})\!+\!\eta_{m_{1}},\\ \vdots\\ \!\Delta_{N}(t)\!=\!\Delta_{N\!-\!1}(t\!-\!m_{N\!-\!1})\!+\!\eta_{m_{N\!-\!1}},\\ \!\Delta_{N\!+\!1}(t)\!=\!\Delta_{N}(t\!-\!m_{N})\!+\!\eta_{m_{N}},\end{cases}\hskip-10.0pt\!\Rightarrow\!\Delta_{N\!+\!1}(t)\!=\!\Delta_{1}(t\!-\!\!\underbrace{\sum_{i=1}^{N}m_{i}}_{=\tau_{N}})\!+\!\!\underbrace{\sum_{i=1}^{N}\eta_{m_{i}}}_{=\beta_{N}}\!.$
@@
$\!\overset{(a)}{=}\!{\sum_{\tau=N}^{\infty}\!\mathbb{P}(\tau_{N}\!=\!\tau)\mathbb{E}\left[\Delta_{1}(t\!-\!\tau)\right]}\!+\!p_{g}\!\sum_{i=1}^{N}\frac{1}{\rho_{i}},$
@@
Equality $(b)$ follows directly from the relation $\lim_{t\rightarrow\infty}\mathbb{E}[\Delta_{1}(t-\tau)]=\lim_{t\rightarrow\infty}\mathbb{E}[\Delta_{1}(t)]$, which holds for an ergodic and integrable dtmc $\Delta_{1}(t)$, where $\mathbb{E}\big[|\Delta_{1}(t)|\big]<\infty$ [31, Sec. 1.10]. This condition is satisfied by the dtmcs presented in Section IV under all three policies, for which the steady-state distribution is stationary and therefore time-invariant. ∎